OpenStudy (anonymous):

Factor completely: 8m^3 - 12m^2 + 4m Answer 4 (2m^3 - 3m^2 + m) 4m (2m - 1)(m - 1) 2m (4m^2 - 6m + 2) 4m (2m^2 - 3m)

OpenStudy (anonymous):

4m(2m-1)(m-1) 8m^3 -12m^2+4m

OpenStudy (ajprincess):

4m is common to all three terms. so take it out. when u do so u will get 4m(2m^2-3m+1) =4m(2m^2-2m-m+1) =4m(2m(m-1)-(m-1) =4m(2m-1)(m-1)

OpenStudy (ajprincess):

Is that clear?

OpenStudy (anonymous):

yea